OpenStudy (anonymous):

how do i express the number 29.73213749 in a radical form?

OpenStudy (amistre64):

why would you want to?

OpenStudy (amistre64):

... did you get that on your calculator and need to match it up to an option by chance?

OpenStudy (amistre64):

\[a = x^{1/n}\] \[a^n = x\]

OpenStudy (anonymous):

I got that number on my calculator after finding the square root of 884 and i was ask to put it in radical form if its not an integer

OpenStudy (amistre64):

you might just want to work the radical on paper then ...

OpenStudy (anonymous):

but i dont know how to do radicals

OpenStudy (amistre64):

its a sqrt ... just start halving it until you get to an odd value

OpenStudy (amistre64):

884 2 * 442 2 * 221 etc ... what are any factors of 221?

OpenStudy (anonymous):

13 and 17

OpenStudy (amistre64):

and those are prime so it looks like 884 = 2*2*17*13 sqrt(884) = sqrt(2*2*17*13) anything that is doubled can be pulled out sqrt(2*2*17*13) = 2 sqrt(17*13) so id say its simply 2 sqrt(221)

OpenStudy (anonymous):

Thank you !

OpenStudy (amistre64):

your welcome, good luck :)